OpenStudy (anonymous):

Please Help! The point (a,b) is on the graph of f and the line tangent to the graph at (a,b) passes through the point (0,-8) which is not on the graph of f. Find the values of a and b.

OpenStudy (anonymous):

f(x)= x^3 - x^2 - 4x + 4

OpenStudy (anonymous):

First of all, remember that the derivative of that equation is the line tangent to it.

OpenStudy (anonymous):

f '(x)=3x^2 -2x - 4

OpenStudy (anonymous):

now what do I do?

OpenStudy (anonymous):

Hmm, I guess derive again to get the tangent line, since that's still a curve.

OpenStudy (anonymous):

Then find a line that's consistent with (a, b) given.

OpenStudy (anonymous):

so find the slope of the tangent line in terms of a and b and plug in 0 and -8?

OpenStudy (anonymous):

hold on, doing some research. I think I may have led you astray - I think the right thing is probably to just take the first derivative and then plug that in as the slope of a line in point-slope form. Where point is (x, f(x)), and the slope is f'(x).

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

right, where x1 is x, y1 is f(x), and m is f'(x)

OpenStudy (anonymous):

no x1 and y1 is the point

OpenStudy (anonymous):

To keep the x's and y's from getting mixed up, use a different x and y for the point-slope equation -- e.g. xx and yy yy- f(x) = f'(x)(xx - x)

OpenStudy (anonymous):

Then plug in the point (0, -8) you're given as xx and yy, and solve for x.

OpenStudy (anonymous):

So it would be -8 - f(x) = f'(x)(0 - x), and solve for x.

OpenStudy (anonymous):

I think that should do it.

OpenStudy (anonymous):

did that make sense?

OpenStudy (anonymous):

i think that is wrong for some reason, I have never seen the point slope formula been used like that before

OpenStudy (anonymous):

Neither had I until I looked up how to find the equation of a tangent line to a curve on wikipedia just now

OpenStudy (anonymous):

the answer is a=2 and b=0

OpenStudy (anonymous):

thats not what I necessarily got, but thats what her answer key says

OpenStudy (anonymous):

hold on, let me type my equation into mathematica and see if that's the answer it gives

OpenStudy (anonymous):

yep, it solves to 3 solutions, one of which is real and is x=2

OpenStudy (anonymous):

and f(2) = 0

OpenStudy (anonymous):

i still feel like there was a better way of doing this

OpenStudy (anonymous):

cause I'm sell confused

OpenStudy (anonymous):

still*

OpenStudy (anonymous):

well, to define any line, you need two things: a point on the line, and its slope.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

At any point x on your curve, the tangent line is a line that passes through the curve -- i.e., the point (x, f(x)), and has a slope f'(x). Does that make sense?

OpenStudy (anonymous):

To define a tangent line at x, we need a point and a slope. The slope by definition is f'(x), and a point is (x, f(x)).

OpenStudy (anonymous):

yes the slope of the tan line is the deriv. of f(x) and the point is usually either given or just the x value and you plug in the x value to f(x) to get the y value too

OpenStudy (anonymous):

Right, the slope is the deriv of f(x),and we know one point is (x, f(x)). we have to find the value of X such that the line also passes through (0, -8). so the way I found it is used the point-slope form of the line to define the line, then solved it to see which of those tangent lines is also consistent with a=0, b=-8.

OpenStudy (anonymous):

its like backwards from what I usually do, no?

OpenStudy (anonymous):

they didn't give the "x" value they just said it was a?

OpenStudy (anonymous):

Ah - they used a, b, that's better than xx, yy. So the point-slope form of a line, if we're using a and b instead of x and y, is: (b - y1) = m(a-x1) We know a point on the tangent line is (x, f(x)), that's your x1 and y1. We know another point is (0, 8) so you plug that in for a, b. We also know m = f'(x), so you plug that in. You get a big equation that you solve for x, and when you do it solves to x=2. (and 2 complex numbers)

OpenStudy (anonymous):

k stay here imma solve that

OpenStudy (anonymous):

k

OpenStudy (anonymous):

wait it's 0, -8 not 0,8, sorry, that was a typo.

OpenStudy (anonymous):

i know i gotcha

OpenStudy (anonymous):

(-12 + 2x^3 - x^2)=0

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

imma get three answers for that, and I'm terrible at factoring...

OpenStudy (anonymous):

well you can factor x^2 out of the last 2 terms and get the real 0

OpenStudy (anonymous):

(-12 + 2x^3 - x^2) = (-2 + x) (6 + 3 x + 2 x^2)

OpenStudy (anonymous):

-12 + (x^2)(2x-1)=0

OpenStudy (anonymous):

i guess with the factoring above you get one of the zeroes just by inspection at x=2, and the other two by using the quadratic equation on the second term

OpenStudy (anonymous):

sorry ya lost me

OpenStudy (anonymous):

should i do p?'s over q?'s ???

OpenStudy (anonymous):

well if you get to this: (-12 + 2x^3 - x^2) = (-2 + x) (6 + 3 x + 2 x^2) = 0 .. then you know it's equal to 0 if either -2+x = 0, or x^2 + 3x + 6 = 0,. so the -2 + x = 0 you can solve easily, and the other two roots you can get by plugging into the quadratic equation.

OpenStudy (anonymous):

no see i now how to solve from x after you factor, but, how do you get from (-12 + 2x^3 - x^2) to (-2 + x) (6 + 3 x + 2 x^2) = 0

OpenStudy (anonymous):

without like a calc

OpenStudy (anonymous):

haha, i don't actually know, i suck at factoring too to be honest. I used a calculator.

OpenStudy (anonymous):

>_<

OpenStudy (anonymous):

i can't use one UGH

OpenStudy (anonymous):

my final is tomorrow and friday

OpenStudy (anonymous):

and this is part c of question 1 of my free response

OpenStudy (anonymous):

this is a practice question or your actual exam?

OpenStudy (anonymous):

this is one of the practice questions from our AP CALC AB review packet problems

OpenStudy (anonymous):

tomorrow is the "free response" section

OpenStudy (anonymous):

ah. hmm... well, let me try and find an online factoring tutorial ;-)

OpenStudy (anonymous):

so I'm doing the free response questions in our packet

OpenStudy (anonymous):

lol factoring isnt even what our thing is on

OpenStudy (anonymous):

its like limits, derivatives, optimization, theorems, related rates

OpenStudy (anonymous):

yeah, pretty silly that's where we're stuck

OpenStudy (anonymous):

I forgot factoring 20 years ago because the computer usually just does it for you anyway

OpenStudy (anonymous):

I hate high school, I can't wait to get the heck out of here and move on to college

OpenStudy (anonymous):

yeah, math education doesn't get much better there. too much focus on the mechanics (stupid stuff like factoring) and not enough focus on how to model problems (the interesting part of what we did on this problem)

OpenStudy (anonymous):

well hopefully i will be able to test out of it with a good AP score this may

OpenStudy (anonymous):

not to discourage you (you sound like a smart guy or girl) but, i tested out of early calc with my calc AB AP exam, and found every other math class I took was the same, way too focused on how to do the computations and not enough on understanding the concepts

OpenStudy (anonymous):

I didn't really start to enjoy math until I started using programs like mathematica and ocatave that would do all the computations for me, freeing me to actually *understand* the math

OpenStudy (anonymous):

not that this helps you for tomorrow

OpenStudy (anonymous):

well i will be studying business so my math stuff will be in a bit different field, but thats okay i love all forms of math (WHEN IM NOT PRESSURED and I GET TIME TO LEARN), and i have been on my hs's math team for 4 years.

OpenStudy (anonymous):

cool! being good at math will set you apart in business. what college?

OpenStudy (anonymous):

can i tell u in a private chat?

OpenStudy (anonymous):

is that possible on here??

OpenStudy (anonymous):

no .. that's ok, not important

OpenStudy (anonymous):

well to make a long story short

OpenStudy (anonymous):

the common app's don't tell u till like xmas time so thats coming up soon, but to two other schools I have applied to that were not common app, I got in to, with money, direct admittance to their business schools and their Honors business programs, those schools being Indiana U and the U. of Texas. My common apps are Boston College, Villanova, Notre Dame, U. So. Cali (USC), U. of San Diego, Santa Clara U, Maryland U,

OpenStudy (anonymous):

and Upenn (wharton)

OpenStudy (anonymous):

wow, all excellent schools

OpenStudy (anonymous):

haha ya very scattered but I'm intentionally trying to go as far away from home as possible

OpenStudy (anonymous):

haha, you dont like home?

OpenStudy (anonymous):

do you know what stir crazy is?

OpenStudy (anonymous):

that's me....everyday

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

what part of the US are you in?

OpenStudy (anonymous):

Chicago is one of the best cities in the world, but I need to branch out. I seriously want to see the world out there, and it seems silly to just stay in one part of the world for most of your life.

OpenStudy (anonymous):

totally agree. travel makes people 10x more interesting, 10x more understanding of cultures not their own, 10x more empathic

OpenStudy (anonymous):

the least tolerant people I know invariably are the least traveled

OpenStudy (anonymous):

I did this camp at georgetown 2 summers ago for 5 weeks. i meet people from everywhere in the world, in particular, 6 girls from Puerto Rico, who turned out to be my best friends i will ever have. I visited them last winter and it was just so cool. Im a guy by the way, so when I came back to school most of my friends where shocked to see me hanging out with all these people, especially the good looking P.R. girls haha

OpenStudy (anonymous):

but if i never ventured, I would of never met them, and that would be a shame

OpenStudy (anonymous):

haha, nice

OpenStudy (anonymous):

and camp seems lame, it was a college enrichment program, u take 2 college courses and for fun do w/e u want around DC

OpenStudy (anonymous):

that's where i grew up

OpenStudy (anonymous):

Gtown is very nice!

OpenStudy (anonymous):

lol loved waking up everyday near M street

OpenStudy (anonymous):

yep, i lived in dupont circle for a couple of years, baltimore for 5 (in college), then israel, then LA, then seattle

OpenStudy (anonymous):

wow all over haha thats the way !